OpenStudy (anonymous):

Find P(A and B) P(A) = 3/5, P(B given A) = 5/8

OpenStudy (anonymous):

P(B/A)=P(AandB)/P(A) P(AndB)=P(BandA) P(AandB)=P(B/A)P(A) =(5/8)(3/5) = 3/8